OpenStudy (leyla.tas123):

Suppose a satellite orbits the earth at about 13^4 km per hour. How long would it take to complete 24 orbits, which is a distance of about 13^5 km?

OpenStudy (skullpatrol):

any ideas?

OpenStudy (leyla.tas123):

i think i should keep the base and add the exponents. is that correct?

OpenStudy (skullpatrol):

not really, ask yourself how many times does 13^4 go into 13^5?

OpenStudy (leyla.tas123):

im not sure

OpenStudy (skullpatrol):

$$\Huge \dfrac{13^5}{13^4}$$

OpenStudy (skullpatrol):

That^ expression comes from t = d/r

OpenStudy (skullpatrol):

$$\dfrac{13•13•13•13•13}{13•13•13•13}$$

OpenStudy (leyla.tas123):

oh, so i'm supposed to subtract the exponents

OpenStudy (leyla.tas123):

so, the answer is 13^1

OpenStudy (skullpatrol):

Yes. Don't forget the units :-)

OpenStudy (leyla.tas123):

thank you so much!

OpenStudy (skullpatrol):

Thanks for trying to learn.

OpenStudy (skullpatrol):

$$\Huge\dfrac{\cancel{13•13•13•13•}13}{\cancel{13•13•13•13}}$$